OpenStudy (anonymous):

what is the deravitive of 4^x?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

4^x ln(4)

OpenStudy (anonymous):

i didnt see the raise to a power. you need to implicitly differentiate

OpenStudy (anonymous):

4^x=y, ln(y)/ln(4)=x, (d/dx)ln(y)=(d/dx)ln(4)x, (1/y)(dy/dx)=ln(4), dy/dx=yln4, dy/dx=(4^x)ln(4).

OpenStudy (anonymous):

y = 4^x ln y = x ln 4 implicit derivative, (1/y) y' = ln 4 y' = y ln 4 y' = 4^x ln 4